OpenStudy (anonymous):

find an equation of the line containing the given pair of points (1,5) & (3,6)

OpenStudy (anonymous):

the answer i have is m=1/2 how do i put it in slope-intercept form

OpenStudy (anonymous):

plug in one of the point and solve for the y-intercept (b in equation) y = mx + b since u know x y and m plug in 5 = (1/2)1 +b 5 - 1/2 =b 9/2 = b y = (1/2)x + 9/2

OpenStudy (anonymous):

yup or u could use y= mx +c y=6,m=1/2 and x= 3 you would have 6=1/2(3) + c 6-3/2 = c c= 9/2 therefore y=(1/2)x + 9/2

OpenStudy (anonymous):

is d ans x-2y+9=0